OpenStudy (anonymous):

The volume of water in a reservoir decreases 21,000 cubic yards over a 3-month period. What is the average in volume per month? @mayankdevnani @pooja195 @welshfella @Error1603

OpenStudy (faiqraees):

average DECREASE in volume per month?

OpenStudy (anonymous):

Yes @FaiqRaees

OpenStudy (faiqraees):

\[\large\rm Decrease = \frac{decrease~in~volume}{number~of~months}\]

OpenStudy (anonymous):

Plz Help

OpenStudy (anonymous):

Plz @zepdrix

OpenStudy (anonymous):

Plz @Error1603

OpenStudy (anonymous):

I will give a medal!

OpenStudy (error1603):

21000/3=7000 cubic yards/month

OpenStudy (faiqraees):

@Error1603 Please resist from giving direct answers. It is against the code of conduct and does not help anyone learn. Plus it discourages other inspirational teachers to teach properly since the provision of the answer before the solution, makes the questioner to skip the solution part.

OpenStudy (anonymous):

How did you get this I just don't understand the question.

OpenStudy (anonymous):

Plz help explain.

OpenStudy (error1603):

so you lose 21,000 cubic yards in 3 months. how many cubic yards do you use a month?

OpenStudy (anonymous):

OHHHH ok

OpenStudy (anonymous):

So I divide 21,000 by 3?

OpenStudy (error1603):

Yes

OpenStudy (anonymous):

an that give me....

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

7,000

OpenStudy (anonymous):

Like you said

OpenStudy (anonymous):

Thank you so much @Error1603

OpenStudy (error1603):

No problem

OpenStudy (anonymous):

Wait before you leave i have another question i dont understand

OpenStudy (error1603):

ok

OpenStudy (anonymous):

From midnight to 6:00 am, the temperature rose 8 degrees c. at 6:00 am, the temperture was -20 degrees c. What was the temperture at midnight?

OpenStudy (error1603):

Alright lets see

OpenStudy (anonymous):

k

OpenStudy (error1603):

Sorry I don't really get it myself

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

help anyone

OpenStudy (error1603):

@zepdrix

OpenStudy (anonymous):

That guys doesn't help me and i don't know why I think he is mad at me or something

OpenStudy (anonymous):

plz @zepdrix

OpenStudy (error1603):

oh

OpenStudy (error1603):

I'll try to find someone

OpenStudy (anonymous):

thanks

OpenStudy (error1603):

@daisyrainbowmist

OpenStudy (error1603):

@FaiqRaees

OpenStudy (anonymous):

I need help on the second question i asked on here plz help

OpenStudy (faiqraees):

Faiq raees at service

OpenStudy (error1603):

Great

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Its the one about the midnight

OpenStudy (faiqraees):

From midnight to 6:00 am, the temperature rose 8 degrees c. at 6:00 am, the temperture was -20 degrees c. What was the temperture at midnight? At 6 = -20 Change in temperature= +8 Change in temp = Temp at 6 am-Temp at midnight 8 = -20 -x Solve for x

OpenStudy (anonymous):

I still don't understand @FaiqRaees

OpenStudy (faiqraees):

Which part you didnt understand?

OpenStudy (anonymous):

With the 8.

OpenStudy (anonymous):

What do I do with the eight?

OpenStudy (faiqraees):

8=-20-x 8+20 = -x 28=-x x=-28

OpenStudy (anonymous):

At 6 its -20 I understand. But not the +8

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

So you add eight.